Grade 12th passModern Physics

The angle between the vectors i-mj and j+k is π/3 then what is the value of m.

I MOHAMED SHINAZ
8 Years agoGrade 12th pass

1 Answer

Rishabh Gaikwad
8 Years ago
If A=i-mj and B=j+k and theta = pi/3
then apply it to the formula,
                    cos theta = A.B/AB
m comes as sqrt1/3